# ACT: Agentic Classification Tree

### v2

@@ B.2 Qualitative analysis of the questions generated by ACT
In addition to its tree structure, another crucial element ensuring the interpretability of ACT lies in how relevant the generated questions are. For this purpose, we propose to compare these questions with the ones that domain experts would have asked, focusing on the DIAGNO dataset. We survey several relevant medical sources to collect a list of symptoms that are commonly used for tuberculosis diagnosis. These sources include websites of the World Health Organization [^4] https://www.who.int/news-room/fact-sheets/detail/tuberculosis and the British National Health Service [^5] https://www.nhs.uk/conditions/tuberculosis-tb/, and a literature review on tuberculosis diagnosis Storla et al. (2008).
Table 3 shows a side-by-side comparison of the symptoms identified by ACT as positively associated with tuberculosis, and those most commonly identified in the aforementioned medical sources, differentiating the symptoms identified by both ACT and the medical sources (left column), from those that were exclusive (center and right columns). We see that most of the symptoms appear in common. Some of the differences may be explained by ambiguity between some symptoms (e.g. ”severe pain” vs. Chest pain and headaches). Others may come from binary classification setting considered, opposing tuberculosis cases to allergy, potentially resulting in some symptoms (e.g. loss of appetite) not being represented in the dataset.
Overall, this tends to show that ACT is indeed able to identify relevant symptoms to make its predictions, showcasing its utility and reliability.
B.3 Agentic Classification Tree for SPAM Dataset
[…]
Does the text contain promotional
language or marketing terms?
Does this text use excessive exclamation
points, grammatical errors, or
poor writing quality?
SPAM
Does this example primarily aim to
generate immediate engagement
(likes, comments, shares)?
SPAM
HAM
Does this text rely heavily on
emotionally charged language
or sensational claims?
SPAM
Does this example aim to generate
immediate revenue or sales?
SPAM
HAM
yesyesnoyesnonoyesnoyesno
@@ B.3 Agentic Classification Tree for SPAM Dataset
B.4 Agentic Classification Tree for Jailbreak Dataset
[…]
Does this text explicitly instruct an AI to
bypass ethical, legal, or moral constraints?
Does the prompt implicitly or explicitly instruct
the AI to override policy or safety restrictions?
JB
Is this text a direct instruction to generate
harmful, unethical, or illegal content in
a role-playing or creative context?
JB
Not JB
Does the example require imaginative or
persona-driven instructions that guide the
response toward creative or roleplay content?
Does this roleplay scenario explicitly request
or involve role-playing as an AI (e.g., ’I am DAN’,
’I am JailBreak’) that ignores ethical,
legal, or moral guidelines?
JB
Not JB
Is this text a neutral, educational, or
role-playing scenario with fictional characters
in a non-offensive context?
JB
Not JB
yesyesnoyesnonoyesyesnononoyes
@@ B.4 Agentic Classification Tree for Jailbreak Dataset
B.5 Qualitative comparison with TF-IDF + CART
@@
nose $\leq 0.074$
itching $\leq 0.052$
nausea $\leq 0.192$
TB
Not TB
also $\leq 0.294$
Not TB
TB
nasal $\leq 0.238$
constant $\leq 0.289$
Not TB
TB
nose $\leq$ 0.309
TB
Not TB
yesyesyesnononoyesnoyesnoyesnonoyes
@@
[…]
our $\leq 0.041$
your $\leq 0.06$
now $\leq 0.055$
HAM
@@
any $\leq 0.038$
SPAM
HAM
announce $\leq 0.1$
severe $\leq 0.13$
SPAM
HAM
HAM
yesyesyesnonoyesnonoyesyesnono
Appendix: Use of Language Models

### v3

@@ D.1 Qualitative analysis of the questions generated by ACT
In addition to its tree structure, another crucial element ensuring the interpretability of ACT lies in how relevant the generated questions are. For this purpose, we propose to compare these questions with the ones that domain experts would have asked, focusing on the DIAGNO dataset. We survey several relevant medical sources to collect a list of symptoms that are commonly used for tuberculosis diagnosis. These sources include websites of the World Health Organization [^5] https://www.who.int/news-room/fact-sheets/detail/tuberculosis and the British National Health Service [^6] https://www.nhs.uk/conditions/tuberculosis-tb/, and a literature review on tuberculosis diagnosis (storla2008systematic).
Table 5 shows a side-by-side comparison of the symptoms identified by ACT as positively associated with tuberculosis, and those most commonly identified in the aforementioned medical sources, differentiating the symptoms identified by both ACT and the medical sources (left column), from those that were exclusive (center and right columns). We see that most of the symptoms appear in common. Some of the differences may be explained by ambiguity between some symptoms (e.g. ”severe pain” vs. chest pain and headaches). Others may come from the binary classification setting considered, opposing tuberculosis cases to allergy, potentially resulting in some symptoms (e.g. loss of appetite) not being represented in the dataset. Overall, this tends to show that ACT is indeed able to identify relevant symptoms to make its predictions, showcasing its utility and reliability.
D.2 Agentic Classification Tree for SPAM and JAILBREAK Dataset
{forest}
@@ D.2 Agentic Classification Tree for SPAM and JAILBREAK Dataset
{forest}
@@
D.3 Qualitative comparison with TF-IDF + CART
@@
{forest}
@@
{forest}
Appendix E Additional Results and Baseline Implementation details
E.1 Additional Results
To complement the results shown in Table 2 we expand this table in Table 6 and Table 7, which include additional results for the ACT $(d=3,k=20)$ and $(d=4,k=10)$ configurations, as well as strong but not interpretable baselines:
  • TF-IDF + XGBoost: a gradient-boosted tree ensemble trained on TF-IDF representations of the texts (chen2016xgboost). We tune both the maximum tree depth and the number of trees (see Appendix E), and report the best-performing configuration per dataset.
  • Fine-tuned BERT and RoBERTa classifiers: supervised Transformer baselines initialized from pretrained BERT and RoBERTa encoders (devlin2019bert; liu2019roberta) and fine-tuned end-to-end on each dataset for text classification. These models provide strong predictive performance references, but do not yield explicit, human-readable decision rules.
Table 6: Comparison of classification methods across the first three datasets and four LLMs. Results show accuracy and F1-score on the test set (in %) for nine benchmarks, as well as the proposed ACT with varying hyperparameters ($d$: tree depth, $k$: optimization steps per node).
Method Gemma3 4b GPT-4.1 Nano GPT-4.1 Mini Qwen3 4b Avg
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
DIAGNO
CoT (0-shot) 61.5 69.5 63.3 54.8 61.2 51.4 63.2 41.7 62.3 54.4
DSPy (BFSR, 8 demos) 64.0 68.2 68.7 61.2 64.5 59.4 63.3 54.9 64.0 60.9
TextGrad 63.3 56.7 64.5 53.0 65.8 55.6 64.0 60.0 64.4 56.3
TF-IDF + CART (d=3d{=}3) -- -- -- -- -- -- -- -- $78.8$ $78.9$
TF-IDF + XGBoost (d=4,nt=200d{=}4,nt{=}200 ) -- -- -- -- -- -- -- -- $83.0$ $83.5$
BERT -- -- -- -- -- -- -- -- $90.3$ $91.1$
RoBERTa -- -- -- -- -- -- -- -- $89.8$ $90.4$
Con. tag. $65.5$ $63.9$ $59.3$ $58.1$ $75.3$ $78.6$ $58.8$ $64.7$ $64.7$ $66.3$
Rule Fit -- -- -- -- -- -- -- -- $80.2$ $79.3$
ACT (d=3d{=}3, k=10k{=}10) 65.3 66.4 66.8 67.0 77.3 75.2 64.8 61.3 68.6 67.7
ACT (d=3d{=}3, k=20k{=}20) $67.0$ $67.9$ $67.2$ $69.0$ $76.2$ $76.9$ $65.5$ $61.8$ $69.0$ $68.9$
ACT (d=4d{=}4, k=10k{=}10) $66.3$ 71.1 70.5 $70.1$ $74.3$ $70.1$ $66.7$ $66.8$ $69.5$ $69.5$
ACT (d=4d{=}4, k=20k{=}20) 68.3 70.8 70.3 70.8 82.3 81.8 70.3 73.5 72.8 74.2
@@ E.1 Additional Results
CoT (0-shot) 73.3 78.6 95.5 95.3 95.7 95.5 93.2 93.6 89.4 90.8
DSPy (BFSR, 8 demos) 95.0 95.2 98.3 98.3 98.2 98.2 97.2 97.2 97.2 97.2
TextGrad 95.0 94.7 97.3 97.3 96.7 96.6 96.0 96.0 96.3 96.2
TF-IDF + CART (d=5d{=}5) -- -- -- -- -- -- -- -- $92.7$ $93.1$
TF-IDF + XGBoost (d=3,nt=200d{=}3,nt{=}200 ) -- -- -- -- -- -- -- -- $96.3$ $96.3$
BERT -- -- -- -- -- -- -- -- $99.5$ $99.5$
RoBERTa -- -- -- -- -- -- -- -- $99.3$ $99.3$
Con. tag. $83.2$ $85.5$ $64.2$ $45.8$ $82.3$ $85.0$ $62.3$ $39.6$ $73.0$ $64.0$
Rule Fit -- -- -- -- -- -- -- -- $97.0$ $96.9$
ACT (d=2d{=}2, k=5k{=}5) 95.8 95.8 98.7 98.6 98.2 98.2 96.7 96.6 97.4 97.3
ACT (d=2d{=}2, k=10k{=}10) $96.5$ $96.6$ $98.8$ $98.8$ $99.2$ $99.2$ $98.5$ $98.5$ $98.4$ $98.4$
ACT (d=3d{=}3, k=5k{=}5) $98.2$ $98.1$ $97.0$ $96.9$ $99.0$ $99.0$ $98.5$ $98.5$ $98.3$ $98.3$
ACT (d=3d{=}3, k=10k{=}10) 98.5 98.5 99.2 99.2 99.5 99.5 98.8 98.8 99.0 99.0
JAILBREAK
CoT (0-shot) 77.9 82.2 85.5 83.5 93.6 93.3 81.1 77.7 84.5 84.2
DSPy (BFSR, 8 demos) 89.6 89.3 91.6 91.6 94.8 94.8 85.5 83.9 90.4 89.9
TextGrad 91.6 91.3 90.4 90.3 94.8 95.1 88.8 90.0 91.4 91.7
TF-IDF + CART (d=4d{=}4) -- -- -- -- -- -- -- -- $92.8$ $92.4$
TF-IDF + XGBoost (d=2,nt=200d{=}2,nt{=}200 ) -- -- -- -- -- -- -- -- $96.8$ $96.8$
BERT -- -- -- -- -- -- -- -- $98.4$ $98.4$
RoBERTa -- -- -- -- -- -- -- -- $99.6$ $99.6$
Con. tag. $82.7$ $82.2$ $58.6$ $31.8$ $82.7$ $79.8$ $76.3$ $70.1$ $75.1$ $66.0$
Rule Fit -- -- -- -- -- -- -- -- $96.8$ $96.8$
ACT (d=3d{=}3, k=10k{=}10) 82.3 82.7 85.5 85.1 95.2 95.5 82.7 82.6 86.4 86.5
ACT (d=3d{=}3, k=20k{=}20) $84.3$ $86.5$ $91.2$ $90.5$ $96.2$ $96.6$ $87.6$ $86.6$ $89.8$ $90.1$
ACT (d=4d{=}4, k=10k{=}10) $90.0$ $90.1$ $91.6$ $91.0$ $97.2$ $97.2$ $85.1$ $83.9$ $91.0$ $90.6$
ACT (d=4d{=}4, k=20k{=}20) 90.8 90.0 92.0 91.5 98.8 98.8 92.0 92.7 93.4 93.3
Table 7: Comparison of classification methods across the two remaining datasets and four LLMs. Results show accuracy and F1-score on the test set (in %) for nine benchmarks, as well as the proposed ACT with varying hyperparameters ($d$: tree depth, $k$: optimization steps per node).
Method Gemma3 4b GPT-4.1 Nano GPT-4.1 Mini Qwen3 4b Avg
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
BANKCHURN
CoT (0-shot) 47.8 48.1 48.5 25.9 52.8 60.5 50.2 37.3 49.8 43.0
DSPy (BFSR, 8 demos) 52.8 55.6 52.3 42.8 57.2 58.5 54.0 52.2 54.1 52.3
TextGrad 52.0 60.0 53.3 60.3 55.3 61.0 52.8 61.4 53.4 60.7
TF-IDF + CART (d=5d{=}5) -- -- -- -- -- -- -- -- $63.2$ $66.8$
TF-IDF + XGBoost (d=2,nt=200d{=}2,nt{=}200 ) -- -- -- -- -- -- -- -- $56.7$ $48.8$
BERT -- -- -- -- -- -- -- -- $71.7$ $71.6$
RoBERTa -- -- -- -- -- -- -- -- $72.3$ $72.3$
Con. tag. $57.2$ $60.2$ $52.0$ $59.8$ $59.7$ $60.8$ $54.3$ $61.6$ $55.8$ $60.6$
Rule Fit -- -- -- -- -- -- -- -- $63.0$ $59.8$
ACT (d=3d{=}3, k=10k{=}10) 56.3 42.3 58.8 67.4 60.5 64.6 57.5 62.7 58.3 59.3
ACT (d=3d{=}3, k=20k{=}20) $58.0$ $47.8$ 67.0 $67.5$ 69.2 $65.3$ 62.7 63.7 $64.2$ $61.1$
ACT (d=4d{=}4, k=10k{=}10) $61.7$ $65.3$ $61.7$ $60.9$ $69.0$ $65.0$ $60.9$ $60.2$ $63.3$ $62.9$
ACT (d=4d{=}4, k=20k{=}20) 65.2 70.7 65.0 69.1 68.7 68.6 62.0 60.0 65.2 67.1
IMDB
CoT (0-shot) 92.3 91.9 94.2 94.1 95.7 95.6 94.2 94.0 94.1 93.9
DSPy (BFSR, 8 demos) 93.7 93.6 95.5 95.5 96.8 96.8 94.8 94.8 95.2 95.2
TextGrad 93.2 92.9 95.2 95.1 95.8 95.8 94.7 94.5 94.7 94.6
TF-IDF + CART (d=3d{=}3) -- -- -- -- -- -- -- -- $66.8$ $72.8$
TF-IDF + XGBoost (d=2,nt=200d{=}2,nt{=}200 ) -- -- -- -- -- -- -- -- $76.2$ $76.7$
BERT -- -- -- -- -- -- -- -- $90.2$ $90.1$
RoBERTa -- -- -- -- -- -- -- -- $92.7$ $92.7$
Con. tag. $74.8$ $76.9$ $56.2$ $31.7$ $50.0$ $66.7$ $51.3$ $63.8$ $58.1$ $59.8$
Rule Fit -- -- -- -- -- -- -- -- $75.2$ $73.7$
ACT (d=3d{=}3, k=10k{=}10) 93.2 93.2 92.8 92.8 95.7 95.7 93.8 93.9 93.9 93.9
ACT (d=3d{=}3, k=20k{=}20) 93.7 93.6 94.7 94.7 96.3 96.2 95.5 95.5 95.1 95.0
E.2 Tree-Based Baselines
To complement the prompting-based baselines used in the main paper, we report in Table 8 an extended comparison with purely symbolic lexical models. These include TF–IDF + CART with depths ranging from 2 to 5, and TF–IDF + XGBoost with multiple depth and tree-count configurations. These models provide a transparent, non-LLM control group and clarify the performance gap between ACT and traditional supervised baselines across all five datasets. As shown below, ACT remains competitive to these methods while offering semantically grounded, human-readable decision rules.
Table 8: Performance of TF-IDF + CART and TF-IDF + XGBoost across all datasets.
Dataset / Split CART d=2 CART d=3 CART d=4 CART d=5 XGB (2,50) XGB (2,200) XGB (3,50) XGB (3,200) XGB (4,50) XGB (4,200)
DIAGNO – Train 82.0 84.3 86.5 88.8 86.2 94.7 89.2 98.7 93.5 99.8
DIAGNO – Test 77.5 78.8 73.7 76.3 81.0 81.5 81.8 81.7 82.2 83.0
SPAM – Train 87.8 92.3 93.5 94.7 97.8 99.7 98.7 99.8 99.2 100.0
SPAM – Test 85.8 89.3 90.7 92.7 94.3 96.2 93.8 96.3 94.5 96.0
JAILBREAK – Train 88.3 91.1 93.1 94.3 95.6 97.9 96.5 98.5 97.1 99.0
JAILBREAK – Test 86.4 92.4 92.8 91.9 95.2 96.8 95.2 96.4 95.2 96.8
BANKCHURN – Train 59.7 63.0 63.7 66.7 71.5 84.5 79.3 90.0 84.2 93.2
BANKCHURN – Test 62.7 62.8 62.8 63.2 53.2 56.7 54.3 53.7 54.8 55.3
IMDB – Train 68.2 71.0 71.8 76.2 83.5 97.3 89.0 99.3 94.7 99.8
IMDB – Test 66.8 66.8 66.5 66.8 72.5 76.2 73.3 74.8 74.8 75.5
Appendix F Interpretability Statistics for ACT
In Table 9(e) we report two interpretability indicators for ACT across all datasets and ACT configurations evaluated: the average question length (in characters) and the average path length (number of nodes traversed during inference). These values are averaged over the four LLMs used in our experiments. Across datasets, average question lengths range from 78 to 155 characters, reflecting concise node-level prompts, while average path lengths show that predictions typically require only about three nodes, even for depth-4 trees. We also observe a mild trade-off in which datasets with longer questions (e.g., JAILBREAK and SPAM) yield shorter paths, indicating that ACT can consolidate multiple relevant features into a single informative query. For reference, the generic seed question “Based on the provided example, does it belong to the positive class? (yes/no)” has a length of 87 characters.
Table 9: Interpretability statistics across datasets.
ACT (d, k) Avg. Q. Len Avg. Path
(3, 10) 92.00 2.35
(3, 20) 93.15 2.51
(4, 10) 78.05 3.03
(4, 20) 80.53 3.06
(a) DIAGNO
ACT (d, k) Avg. Q. Len Avg. Path
(2, 5) 118.00 1.33
(2, 10) 142.88 1.26
(3, 5) 94.18 1.36
(3, 10) 122.63 1.14
(b) SPAM
ACT (d, k) Avg. Q. Len Avg. Path
(3, 10) 136.00 1.98
(3, 20) 155.00 2.13
(4, 10) 142.00 2.41
(4, 20) 150.33 2.57
(c) JAILBREAK
ACT (d, k) Avg. Q. Len Avg. Path
(3, 10) 90.50 2.92
(3, 20) 89.90 2.93
(4, 10) 82.20 3.10
(4, 20) 89.77 2.96
(d) BANKCHURN
ACT (d, k) Avg. Q. Len Avg. Path
(3, 10) 81.30 1.97
(3, 20) 95.80 2.28
(e) IMDB
Appendix G Additional Ablation Studies and Scaling Analysis
G.1 Ablation on the Number of Feedback Samples $m$
We conduct an ablation study on the number of labeled samples $m$ used to generate semantic feedback during node-level prompt optimization. We evaluate $m\in\{5,10,20,50\}$ on the DIAGNO dataset using GPT-4.1-nano for two ACT configurations: $(d=3,k=10)$ and $(d=4,k=20)$. The results in Tables 10 and 11 show how varying $m$ affects training performance, test metrics, and token usage. As comparison, the results for DSPy are also shown.
Table 10: Effect of varying $m$ for ACT $(d=3,k=10)$ on DIAGNO with GPT-4.1-nano.
m Train Acc $\uparrow$ Test Acc $\uparrow$ Test F1 $\uparrow$ Tokens ($\times 100k$) $\downarrow$
5 69.8 67.0 67.6 52.4
10 71.2 67.5 71.3 53.5
20 71.3 69.5 71.7 54.2
50 71.8 67.2 69.0 55.2
DSPy 71.5 68.7 61.2 122.1
Table 11: Effect of varying $m$ for ACT $(d=4,k=20)$ on DIAGNO with GPT-4.1-nano.
m Train Acc $\uparrow$ Test Acc $\uparrow$ Test F1 $\uparrow$ Tokens ($\times 100k$) $\downarrow$
5 71.7 65.8 65.9 124.7
10 74.2 68.7 70.6 135.0
20 74.5 70.0 71.3 137.6
50 74.8 70.3 70.8 143.8
DSPy 71.5 68.7 61.2 122.1
G.2 Model Scaling and Cost-Normalized Analysis
To assess how model strength affects different prompting strategies, we conducted an additional scaling study using GPT-4.1. We evaluate CoT, DSPy, TextGrad, and two ACT configurations on the DIAGNO and BANKCHURN datasets, as these are the most challenging tasks. As shown in Table 12, stronger models do not consistently improve CoT, DSPy, or TextGrad, whereas ACT benefits from increased model capacity. These results indicate that ACT more effectively leverages the enhanced reasoning abilities of more powerful models.
Table 12: Performance scaling of CoT, DSPy, TextGrad, and ACT when moving from GPT-4.1-nano to GPT-4.1-mini to GPT-4.1. Values shown are test-set accuracies. ACT benefits more consistently from stronger models.
Method GPT-4.1-nano GPT-4.1-mini GPT-4.1
DIAGNO
CoT 63.3 61.2 62.2
DSPy 68.7 64.5 68.2
TextGrad 64.5 65.8 64.8
ACT (3, 10) 66.9 77.3 80.2
ACT (3, 20) 67.2 76.2 81.8
BANKCHURN
CoT 48.5 52.8 57.2
DSPy 52.3 54.0 59.5
TextGrad 53.3 55.3 58.2
ACT (3, 10) 58.8 60.5 69.2
ACT (3, 20) 67.0 69.2 72.5
Figure 3 in the main paper visualizes the accuracy–cost trade-off across methods and model sizes, highlighting how performance scales with increasing computational expense. This section further reports token usage, monetary cost, test accuracy, and cost-normalized accuracy for each method, with Table 13 summarizing these metrics across the GPT-4.1 family on DIAGNO to facilitate comparison of efficiency and scalability.
Table 13 shows that TextGrad achieves the highest cost-normalized accuracy in the Nano and Mini settings, while ACT (3,10) remains competitive—ranking just behind TextGrad in smaller models and performing strongly on GPT-4.1. The (3,20) variant is less cost-efficient due to higher token usage, though it still slightly outperforms DSPy on Mini and GPT-4.1. As shown in Figure 3 in the main paper, ACT’s test accuracy consistently scales with cost across all model sizes, whereas TextGrad and DSPy do not exhibit consistent cost-dependent accuracy gains.
Table 13: Cost normalization and model scaling across the GPT-4.1 family on the DIAGNO dataset.
Methods Model Cost $\downarrow$ Test Acc. $\uparrow$ Total Tokens (x100) $\downarrow$ Cost Norm Acc. $\uparrow$
3, 10 Nano 1.09 66.90 55.18 61.54
3, 20 Nano 2.02 67.20 101.37 33.31
DSPY Nano 1.85 68.70 152.58 37.23
TextGrad Nano 0.80 64.50 48.06 80.93
3, 10 Mini 3.09 77.30 48.08 24.98
3, 20 Mini 5.58 76.20 87.58 13.65
DSPY Mini 7.92 64.50 161.28 8.15
TextGrad Mini 2.30 65.80 36.37 28.60
3, 10 GPT-4.1 19.08 80.20 53.71 4.20
3, 20 GPT-4.1 36.11 81.80 102.46 2.26
DSPY GPT-4.1 37.10 68.20 153.89 1.84
TextGrad GPT-4.1 17.76 64.80 49.08 3.65
Appendix H Use of Large Language Models